Calculate the discriminant b2- 4ac to determine the number and type of solutions to x2+ 4x + 4= 0

1
Expert's answer
2021-02-24T07:57:04-0500

Solution

Let’s compare given equation with standard form of quadratic equation ax2+ bx + c= 0.

Here a = 1, b = 4, c = 4.

So b2 – 4ac = 42 – 4*1*4 = 16 – 16 = 0.

Therefore given equation has a double root x1,2 = -2.